OpenStudy (anonymous):

14. mrs. Quinn is buying square tiles to cover her kitchen floor. The floor measures 9 feet by 12 feet. Each side of a tile is 4 inches. if Mrs.Quinn buys 950 tiles, are there enough tiles to cover the floor?

OpenStudy (anonymous):

When I first read this I thought it said square tires...that made me laugh so: couple ways to look at this problem - can convert the square feet to square inches by going (9*12)*(12*12) then divide by the 4*4 tiles = 972 tiles or if you have 9 feet across each foot across would need 3 4inch tiles so 9*3 = 27 then to cover the other dimension 12 *3 =36 so need 36 rows of 27 tiles to cover 36*27=972 or divide the number of tiles by the number of tiles across to get the number down 950/27 = 35....so not quite enough to cover